OpenStudy (anonymous):

Which expresses 1050 : 2000 as a decimal? (Points : 1) 0.525 0.38 0.075 0.05

OpenStudy (anonymous):

?

OpenStudy (anonymous):

36:32

OpenStudy (anonymous):

...Not the time to Ask for ?'s...

OpenStudy (campbell_st):

use the unitary method start my recognizing 136 = 2l + 2w so 68 = l + w Step 1: find the total number of parts in the ratio 9 + 8 = ? Step 2: find the value of 1 part of the ratio 1 part = 68/? ? is the answer from step 1 Step 3: multiply the value of 1 part by the ratio values length = 9 x 1 part width = 8 x 1 part hope it helps

OpenStudy (anonymous):

Really? well then...

OpenStudy (anonymous):

Which expresses 216 : 270 in its simplest form? (Points : 1) 10 : 8 8 : 10 5 : 4 4 : 5

OpenStudy (anonymous):

4:5

OpenStudy (campbell_st):

@Immabee55 the aim of Open Study is to assist learning and avoid providing answers

OpenStudy (anonymous):

its simple what of the 2 numbers in the answers is equal for instance 216/4: 270/5

OpenStudy (anonymous):

well i was getting to that at least i give the correct answer and not just say a bunch of mathematical terms people cant understand

OpenStudy (anonymous):

and they both equal 54:54 making it that, post more questions and @immabee55 my name in them i will gladly help :)